OpenStudy (anonymous):

Find the local minimum, maximum, and saddle points of the function g(x,y) = (x2 + y) e^y/2 .

OpenStudy (astrophysics):

Is it \[g(x,y) = \frac{ (x^2+y)e^y }{ 2 }\]

OpenStudy (anonymous):

no just e^(y/2)

OpenStudy (anonymous):

no y/2 is the exponent of e

OpenStudy (astrophysics):

\[g(x,y)= (x^2+y)e^{y/2}\]

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

We have to do the second derivative test, you know it?

OpenStudy (astrophysics):

Lets make it easier on ourselves and expand it \[g(x,y) = x^2e^{y/2}+ye^{y/2}\]

OpenStudy (anonymous):

ok

OpenStudy (astrophysics):

So find all the derivatives, find \[\huge f_x,f_y, f_{xx}, f_{yy}\]

OpenStudy (astrophysics):

Essentially the second derivative rule says, let \[D = D(a,b) = f_{xx}(a,b)f_{yy}(a,b)-[f_{xy}(a,b)]^2\] and if D>0 and fxx(a,b) > 0, then f(a,b) is a local min, if D>0, fxx(a,b)<0 then f(a,b) is a local maximum, and if D<0 then f(a,b) then it is a saddle point as it is neither max/ min.

OpenStudy (astrophysics):

But, lets just go over this and you will see, so find the partial derivatives as I've asked you to

OpenStudy (anonymous):

fx= 2xe^(y/2)

OpenStudy (anonymous):

fy=e^(y/2)(x^2 +y /2) + 1

OpenStudy (anonymous):

fxx=2e^(y/2)

OpenStudy (anonymous):

fyy=1/2 e^(y/2) +1/2 e^(y/2) (x^2 +y +2/2)

OpenStudy (anonymous):

@1069592ppl

OpenStudy (astrophysics):

\[\large g_x = 2xe^{y/2} ~~~ \checkmark\] \[\large g_{xx} = 2e^{y/2}~~~\checkmark\] That looks good, now let me check your partial respect to y.

OpenStudy (astrophysics):

\[\large g_{y} = \frac{ 1 }{ 2 }e^{y/2} (x^2+y+2)\] this should be function with respect to y

OpenStudy (astrophysics):

Then \[\large g_{yy} = \frac{ 1 }{ 4 }e^{y/2}(x^2+y+4)\]

OpenStudy (anonymous):

ok thanks

OpenStudy (astrophysics):

Now since we have all the partial derivatives, now we need to see where \[\large g_x = 0~~~\text{and}~~~g_y = 0\] So lets set it up as such \[\large g_x=2xe^{y/2} =0\] and \[g_y = \frac{ 1 }{ 2 }e^{y/2}(x^2+y+2) = 0\] we set them equal to 0, note we are looking for the conditions now which satisfy both of the equations.

OpenStudy (astrophysics):

These are long problems..I know haha.

OpenStudy (astrophysics):

So for gx we are solving for y

OpenStudy (astrophysics):

\[2xe^{y/2}=0\]

OpenStudy (anonymous):

xey/2=0

OpenStudy (anonymous):

how do we solve for y

OpenStudy (astrophysics):

Mhm sec, I am getting - infinity for that

OpenStudy (astrophysics):

Oh oops, solve for x haha.

OpenStudy (astrophysics):

That makes more sense, and should be a lot more easier :)

OpenStudy (anonymous):

so x=0

OpenStudy (astrophysics):

Right, now for gy solve for y

OpenStudy (astrophysics):

\[\large g_y = \frac{ 1 }{ 2 }e^{y/2}(x^2+y+2) = 0\]

OpenStudy (astrophysics):

So the reason we're doing this is actually to find the critical points, this is the most tedious part, but after it will get better.

OpenStudy (anonymous):

i don't know how to solve it because of e^(y/2)

OpenStudy (astrophysics):

Get rid of it :P